# Breast Histopathology Image Retrieval by Attention-based Adversarially Regularized Variational Graph Autoencoder with Contrastive Learning-Based Feature Extraction

### v2

@@ Abstract
Breast cancer is the most common cancer type in women worldwide. Early detection and appropriate treatment can significantly reduce its impact. While histopathology examinations play a vital role in rapid and accurate diagnosis, they often require a substantial workforce and experienced medical experts for proper recognition and cancer grading. Automated image retrieval systems have the potential to assist pathologists in identifying cancerous tissues, thereby accelerating the diagnostic process. Nevertheless, due to considerable variability among the tissue and cell patterns in histological images, proposing an accurate image retrieval model is very challenging.
This work introduces a novel attention-based adversarially regularized variational graph autoencoder model for breast histological image retrieval. Additionally, we incorporated cluster-guided contrastive learning as the graph feature extractor to boost the retrieval performance. We evaluated the performance of the proposed model on two publicly available datasets of breast cancer histological images and achieved superior or very competitive retrieval performance, with average mAP scores of 96.5% for the BreakHis dataset and 94.7% for the BACH dataset, and mVP scores of 91.9% and 91.3%, respectively.
@@
Computational pathology , Graph Neural Network , Graph Construction , Adversarial Training , Variational Graph Autoencoder , Attention Module , Image Retrieval
@@ 1 Introduction
Breast cancer is a global health concern for women and has led to considerable mortality worldwide [1]. It is the most commonly diagnosed cancer and represents around 25% of all cancers diagnoses in women [2]. According to 2018 statistics, breast cancer accounted for 15% of all cancer-related deaths and has increased by 24% in recent years [3]. While breast cancer causes many deaths worldwide, early detection and diagnosis followed by appropriate treatment can help to reduce its impact [4]. The need for early detection of breast cancer has led to new solutions, such as computer-aided diagnosis systems based on histological images [5]. This technology is a valuable tool to improve breast cancer diagnosis and potentially can save lives [6]. However, proper diagnosis based on histological images often requires a substantial workforce and experienced medical experts. To overcome this challenge, pathologists can benefit from referring to archived databases for complex cases to find similar examples with verified findings to minimize diagnostic errors.
Content-based image retrieval (CBIR) is an approach to identify visually similar images in a database and can be used for various applications [7, 8]. For example, pathologists reviewing tissue sections for breast cancer diagnosis may benefit from searching through similar labeled histopathological image data to improve their diagnostic accuracy in distinguishing cancerous from non-cancerous tissue. CBIR techniques have also been used for various non-medical applications such as ”image search” for general and natural images, visual search for retail products, and multiple tools for recognizing faces and artworks [9]. Prior machine learning-based CBIR systems often employed models adapted to specific applications, requiring the collection of labeled data for each particular use, which presented a significant implementation challenge. Additionally, the term ”similarity” in these approaches could have different interpretations depending on the specific applications [10]. For instance, two histological images might be considered similar because they originate from the same organ, are classified as the same type of cancer, share similar staining, or represent comparable histological features.
Content-based medical image retrieval (CBMIR), a subset of CBIR, is specifically adapted to medical images and has been used in the context of radiology and histology in former studies [9, 10, 11]. CBMIR for histological images, not only aids histopathology in grading new tissue samples but also assists in analyzing patterns within similar tissues from previous patients [11, 5]. CBMIR offers distinct advantages over classification tasks. For instance, in classifying prostate cancer, a classification model provides class labels for each input. In contrast, a CBMIR system produces the top-$k$ most similar images based on a query image, where $k$ represents the number of desired results. Upon availability, it can also provide other relevant information such as clinical meta-data or diagnosis reports [12]. Therefore, CBMIR provides pathologists and practitioners with more comprehensive information about diverse tissue patterns and enables them to identify the most similar patterns within different tissues, which could potentially minimize intra-and inter-observer variability [9, 13]. Another advantage of CBMIR is that it inherently includes a clinician-in-the-loop, which is often required for artificial intelligence (AI) applications in healthcare. Furthermore, a pathologist using CBMIR in the diagnostic process is required to come to their own diagnostic conclusion after incorporating additional information provided by CBMIR systems rather than replacing the pathologist’s diagnosis with an AI-based classifier. Unlike histological images with their reports written by pathologists that are presented by CBMIR systems, a classifier is not necessarily easily interpretable and - even if presented to a clinician as a second opinion - may lead to automation bias [14, 15]. From a machine learning perspective, CBMIR is typically performed using unsupervised techniques, while classifiers are usually trained with supervised methods that require labeled datasets.
Histopathology image retrieval is a very challenging task due to the substantial tissue and cell pattern variability [10]. Various semi- and fully-automatic approaches have been proposed for histopathology image retrieval [9]. Among these, graph neural networks (GNN) have gained considerable attention in recent years. Considering the grid data structure in images and the tremendous developments in convolutional neural networks (CNNs), many graph convolutional neural network (GCNN)-based architectures have also been extensively proposed for a wide range of computer vision tasks, including image retrieval [16]. In GCNN, input images can be represented as the graph nodes, and their features are usually extracted by pre-trained CNNs such as Visual Geometry Group (VGG) [17] or EfficientNet [18] to construct the graph. However, the exploited pre-trained models are typically trained on natural images, and their extracted features may not be optimally suited for medical data (e.g., histological images) [19, 20].
@@
While supervised learning is usually used to train GCNNs, some GCNN variants, such as variational graph autoencoder (VGAE) [21], can be used for unsupervised learning tasks. In VGAE, the encoder network maps input data to a lower-dimensional latent space, and the decoder network maps latent space vectors back to the original input data. Unlike conventional graph autoencoders (GAE), VGAE utilizes probabilistic encoding that maps input data to a probability distribution over the latent space. VGAE performance can be further improved by incorporating other techniques, such as adversarial training or adding an attention mechanism to the model [21]. In adversarially regularized variational graph autoencoder (ARVGA) [22], the encoder is forced to generate embedding vectors resembling the data from a prior distribution through adversarial training. Incorporating an attention mechanism into CNNs [23] or GNNs [24] is another technique that has been exploited. These techniques have also been jointly applied for specific applications such as graph link prediction [25], graph clustering visualization [25], and node classification in partially labeled graphs [26].
@@
In this study, for the first time, we adapted and exploited attention-based adversarially regularized variational graph autoencoders (A-ARVGAE) for the histological breast image retrieval task. Moreover, unlike previous studies that use conventional pre-trained CNNs for feature extraction to build the graph, we used the recently developed clustering-guided contrastive learning (CCL) feature extractor [27], which is explicitly designed for histopathological images. We evaluated the performance of our proposed approach on two binary and multi-class histopathological breast image datasets, namely breast cancer histopathological database (BreakHis) [28] and breast cancer histology images dataset (BACH [29]), and achieved superior or very competitive performance compared to other state-of-the-art models. The main contribution of this research can be summarized as follows:
@@
    We adapted and exploited one of the state-of-the-art GNN models, which incorporated both attention mechanism and adversarial training for breast cancer histology image retrieval task.
@@
    Instead of using conventional pre-trained CNNs as the node feature extractor, we made use of CLL, which is explicitly designed for histological images.
@@
    Through extensive experiments, we show the excellent performance of our approach for both utilized GNN and feature extraction on two histopathological breast image datasets, namely BreakHis and BACH, and achieved superior or very competitive performance compared to other approaches.
The rest of this paper is organized as follows: In Section 2, we present a background on methods for medical image retrieval and graph representation learning in histopathology. Section 3 describes our methodology, which contains all steps for the proposed framework. Section 4 consists of the experiments and evaluations with an analysis of the results. Finally, in Section 5, we conclude the paper with suggestions for future works.
@@ 2 Related Works
This section contains a literature review for medical image retrieval and graph representation learning in histopathology, considering their relevance to our work.
@@ 2.1 Medical Image Retrieval
Various methods have been proposed in the literature for medical image retrieval. Text-based image retrieval systems were among the first attempts to find similar images in a given database. However, these methods rely heavily on expert annotation and proper keyword selection, which makes them generally considered ineffective and time-consuming [30].
@@
CBMIR models, on the other hand, rely on the content (features) of the images instead of their given labels. Consequently, they can be used in an unsupervised manner without the need for manual labeling. Conventional CBMIR models consist of image preprocessing, hand-crafted or automatic feature extraction, and feature similarity measurements using different distance-based methods. To enhance interpretability, some approaches propose reducing the disparity between the extracted features and the high-level semantic meanings of the images by incorporating expert feedback or medical metadata into the retrieval model [30, 31].
@@
Commonly used preprocessing methods include noise reduction, contrast enhancement, patch generation (e.g., from whole slide images), and normalization (e.g., stain normalization in histological images) [32, 33]. However, many novel machine learning (ML)- and deep learning (DL)-based feature extractors do not require any preprocessing steps.
For feature extraction, both manually designed and DL-based automatic feature extractors have been utilized in former studies. Hand-crafted features are employed to depict the morphological and textural characteristics of images. These features include color histogram-based features, histogram of oriented gradients (HOG) descriptor, local statistical features, global image structure tensor (GIST) features, Gabor features, wavelet-based features, intensity-based features, textural features, gray level co-occurrence matrix (GLCM) features, and scale-invariant feature transform (SIFT) descriptors [30, 27]. It should be noted that experts usually need to set up specific transformation functions to derive these features. However, more recently, the rapidly evolving field of DL has made it easier to automatically extract high-level features from the images [10].
@@
Many pre-trained DL-based models, such as VGG19, ResNet, MobileNetV2, Xception, EfficientNet, and DenseNet121, have been used as feature extractors for various computer vision tasks such as medical image classification or medical image retrieval [34, 35]. Hand-crafted or automatically extracted features of a query image are then compared to the image features in a database to find similar images. For this comparison, various similarity measurements such as distance-based retrieval (SDR), customized query approach (CQA), or nearest-neighbor matching have been used [36]. In the context of histopathology, Yottixel [37] and FISH [13] methods select the most distinctive patches to represent whole slide images (WSIs) through $K$-means clustering based on RGB histogram and spatial coordinate features. Yottixel employs a pre-trained DenseNet121, trained on the ImageNet dataset [38], to extract patch embeddings, which are then transformed into binary barcodes. FISH trains a model based on an autoencoder on the cancer genome atlas (TCGA) dataset to encode an integer index for each patch during retrieval. However, it is worth mentioning that using a feature extractor trained only on natural images might be sub-optimal due to the fundamental differences between natural and histopathological images. In contrast, CCL is a newly developed model specifically designed for histological images. CCL pretraining was performed with around 15 million image patches extracted from more than 32,000 WSIs.
The CCL learning method utilizes large-scale unlabeled histopathological images for universal feature extraction that can be directly applied for various WSI analysis tasks such as retrieval. In our study, we utilized CCL as a robust and relevant feature extractor to build the graph.
@@ 2.2 Graph Representation Learning in Histopathology
While CNNs have demonstrated impressive performance in medical image analysis, they often have trouble at capturing contextual information from neighboring areas, especially with a limited field of view. In contrast to CNNs, GNNs are better at preserving complex correlations among neighboring elements during the learning process. GNNs have been successfully applied to various tasks, including tumor/object localization, classification, survival prediction, and image retrieval [39].
In the field of computational pathology, several GNN models, such as cell graphs, tissue graphs, cell-tissue graphs, and patch graphs, have been proposed for tasks like node classification (e.g., nuclei classification) or graph classification (e.g., histologic image patch categorization) [40, 41]. However, these applications have primarily been applied in a supervised manner and often need advanced preprocessing. Using other models like HoVerNet [42] or DDU-Net [43] to generate nuclei segmentation masks as input for the graph-based model is an example of such preprocessing. Regarding clinical applications, GNNs have been used in different cancer types analysis, including breast cancer, colorectal cancer, and prostate cancer, among others [41].
To build a graph, one needs to define nodes, node embeddings, and edges. In the context of histological images, nodes could represent nuclei, image patches, or tissue regions. Node embeddings are features of the nodes that can be extracted either manually or automatically using pre-trained models or other techniques. The edges depict connections between nodes, typically determined by predefined criteria such as $k$-nearest neighbors or distances between nodes. Once the graph is built, various GNN or GCN models can be applied, depending on the specific applications.
A GCN contains several key components, including an aggregation layer to define how features from neighbor nodes are aggregated, an activation layer to incorporate non-linearity into the model, graph pooling to reduce graph size (if needed), and graph convolutional operations to update node representations. It’s worth noting that other variants, such as GraphSAGE [44], GAT [45], or ChebNet [46] have been introduced in the literature as novel propagation rules to enhance performance across different applications. The output of a GNN typically depends on the application. For graph classification, a multilayer perceptron (MLP) can be added to the graph. In the case of graph reconstruction for unsupervised tasks, a linear projection decoder can be added to the model. As highlighted in a recent review [39], most GNN-based works in computational pathology are related to supervised classification tasks. Only a few studies have investigated histological image retrieval [39, 27, 47, 12]. However, even in these studies, the potential of attention mechanisms and adversarial training have not been explored. In our research, for the first time, we adapted and made use of both attention mechanism and adversarial training in variational graph autoencoders for the histological image retrieval task.
@@ 3 Methodology
The general workflow of our proposed approach for histological image retrieval is depicted in Figure 1. In the following, we discuss the different parts of the workflow in detail.
Figure 1: The general workflow of the proposed breast histological image retrieval model involves distinct processes for training and testing. A-ARVGAE: Attention-based adversarially regularized variational graph autoencoders.
@@ 3.1 Dataset
The dataset refers to a collection of histological image patches used to train and test the model. Each image has certain features and characteristics related to a specific class or histopathological grade. We use two breast histopathology image datasets in this study, namely the BreakHis dataset [28] and the BACH dataset [29], which are subsequently detailed. Example images from the datasets are shown in Figure 2.
@@
    BreakHis: This dataset contains hematoxylin and eosin (H&E)-stained breast histology microscopy images captured at various magnifications (40×40$\times$40 ×, 100×100$\times$100 ×, 200×200$\times$200 ×, 400×400$\times$400 ×) and is categorized into two primary classes, including benign and malignant classes. The dataset contains 7,909 images including 2,480 benign and 5,429 malignant samples with a size of $700\times 460$ pixels in Portable Network Graphics (PNG) format.
@@ 3.2 Feature Extraction
The first step in our proposed retrieval model is the extraction of features for all images in the dataset. Instead of using hand-crafted features or conventional pre-trained CNNs as feature extractors, we utilize CCL as the main feature extractor. Unlike other conventional pre-trained CNNs, CCL is trained with histological images. We hypothesize that using such a pre-trained feature extractor could lead to better descriptive features for the task of histological image retrieval.
We also use a set of standard baseline models as feature extractors and compare their retrieval performance with the CCL feature extractor. For this, we employ well-known pre-trained models, namely VGG19 [17], MobileNetV2 [48], DenseNet121 [49], Xception [50], InceptionResNetV2 [51], NASNetLarge [52], and EfficientNetV2M [18]. The details of these baseline pre-trained models are presented in Table 1. It is worth mentioning that all these pre-trained models are initially trained on natural images using the ImageNet dataset [38], but the CCL feature extractor used in this study is pre-trained with histological image patches. The results of these comparisons are reported in Section 4 (Table 2 and Table 3).
Table 1: Details of baseline feature extractors. All these pre-trained models are initially trained on natural images using the ImageNet dataset [38].
Feature Extraction Model Size (MB) Parameters Depth
VGG19 549 143.7 M 19
MobileNetV2 14 3.5 M 105
DenseNet121 33 8.1 M 242
Xception 88 22.9 M 81
InceptionResNetV2 215 55.9 M 449
NASNetLarge 343 88.9 M 533
EfficientNetV2M 220 54.4 M -

### v3

# Leveraging Medical Foundation Model Features in Graph Neural Network-Based Retrieval of Breast Histopathology Images

@@ Abstract
Breast cancer is the most common cancer type in women worldwide. Early detection and appropriate treatment can significantly reduce its impact. While histopathology examinations play a vital role in rapid and accurate diagnosis, they often require experienced medical experts for proper recognition and cancer grading. Automated image retrieval systems have the potential to assist pathologists in identifying cancerous tissues, thereby accelerating the diagnostic process. Nevertheless, proposing an accurate image retrieval model is challenging due to considerable variability among the tissue and cell patterns in histological images.
In this work, we leverage the features from foundation models in a novel attention-based adversarially regularized variational graph autoencoder model for breast histological image retrieval. Our results confirm the superior performance of models trained with foundation model features compared to those using pre-trained convolutional neural networks (up to 7.7% and 15.5% for mAP and mMV, respectively), with the pre-trained general-purpose self-supervised model for computational pathology (UNI) delivering the best overall performance. By evaluating two publicly available histology image datasets of breast cancer, our top-performing model, trained with UNI features, achieved average mAP/mMV scores of 96.7%/91.5% and 97.6%/94.2% for the BreakHis and BACH datasets, respectively.
@@
Computational Pathology , Graph Neural Network , Foundation Models , Image Retrieval
@@ 1 Introduction
Breast cancer is a global health concern for women and has led to considerable mortality worldwide [1]. It is the most commonly diagnosed cancer and represents around 25% of all cancers diagnoses in women [2]. According to 2018 statistics, breast cancer accounted for 15% of all cancer-related deaths and has increased by 24% in recent years [3]. While breast cancer causes many deaths worldwide, early detection and diagnosis, followed by appropriate treatment, can help to reduce its impact [4]. The need for early detection of breast cancer has led to new solutions, such as computer-aided diagnosis systems based on histological images [5]. This technology is a valuable tool to improve breast cancer diagnosis and potentially can save lives [6]. However, proper diagnosis based on histological images often requires a substantial workforce and experienced medical experts. Pathologists can benefit from referring to archived databases for complex cases to find similar examples with verified findings to minimize diagnostic errors and overcome this challenge.
Content-based image retrieval (CBIR) is an approach to identifying visually similar images in a database and can be used for various applications [7, 8]. Content-based medical image retrieval (CBMIR), a subset of CBIR, is specifically adapted to medical images and has been used in the context of radiology and histology in former studies [9, 10]. The CBMIR of histological images not only aids histopathology in grading new tissue samples but also assists in analyzing patterns within similar tissues from previous patients [10, 5].
@@
Histopathology image retrieval is challenging due to the substantial tissue and cell pattern variability [11]. Various semi- and fully-automatic approaches have been proposed for histopathology image retrieval [9]. Among these, graph neural networks (GNN) and graph convolutional neural networks (GCNN) have gained considerable attention in recent years [12]. While supervised learning is usually used to train GCNNs, some GCNN variants, such as variational graph autoencoder (VGAE) [13], can be used for unsupervised learning tasks. VGAE performance can be further improved by incorporating other techniques, such as adversarial training or adding an attention mechanism to the model [13]. In an adversarial regularized variational graph autoencoder (ARVGA) [14], the encoder is forced to generate embedding vectors resembling the data from a prior distribution through adversarial training. Incorporating an attention mechanism into convolutional neural networks (CNNs) [15] or GNNs [16] is another technique that has been exploited. These techniques have also been jointly applied for specific applications such as graph link prediction [17], graph clustering visualization [17], and node classification in partially labeled graphs [18].
Generally in recent studies based on GNNs for image retrieval, input images are represented as the graph nodes whose features are usually extracted by pre-trained CNNs to construct the graph. However, the exploited pre-trained CNN models are typically trained on natural images based on the ImageNet dataset [19], and their extracted features may not be optimally suited for medical data (e.g., histological images) [20, 21]. On the other hand, recently, several pre-trained foundation models, trained on millions of data samples in an unsupervised manner, have been developed and applied to a wide range of computer vision tasks.
@@
In this study, we adapted and exploited attention-based adversarially regularized variational graph autoencoders (A-ARVGAE) for the histological breast image retrieval task. Unlike most prior studies that employ conventional pre-trained CNNs for feature extraction to construct the graph, we used medical foundation models, including contrastive language-image pre-training with medical data (BioMedCLIP) [22], general-purpose self-supervised model for computational pathology (UNI) [23], and clustering-guided contrastive learned model (CCL) [24], as the graph node feature extractors. Evaluated on two binary and multi-class histopathological breast image datasets, namely the breast cancer histopathological database (BreakHis) [25] and the breast cancer histology image dataset (BACH) [26], our results demonstrate the superior performance of models trained with foundation model features over those with CNN features, with UNI delivering the best overall retrieval performance. The main contribution of this research can be summarized as follows:
@@
    We adapted and exploited one of the state-of-the-art GNN models, which incorporated both attention mechanism and adversarial training for the histology image retrieval task of breast cancer.
@@
    Instead of using conventional pre-trained CNN features as the node feature extractor, we made use of medical foundation model features.
@@
    Through extensive experiments, we show the excellent performance of our proposed approach for both utilized GNN and feature extraction on two histopathological breast image datasets.
The rest of this paper is organized as follows: In Section 2, we present a brief background on methods for medical image retrieval and graph representation learning in histopathology. Section 3 describes our methodology, which contains all steps for the proposed framework. Section 4 consists of the experiments and evaluations with an analysis of the results. Finally, Section 5 concludes the paper with suggestions for future works.
@@ 2 Related Works
This section contains a brief review of medical image retrieval and graph representation learning in histopathology, considering their relevance to our work.
@@ 2.1 Medical Image Retrieval
Various methods have been proposed in the literature for CBMIR. CBMIR models rely on the content (features) of the images and can be used unsupervised without the need for manual labeling. Conventional CBMIR models consist of image preprocessing, hand-crafted or automatic feature extraction, and feature similarity measurements using different distance-based methods.
@@
Commonly used preprocessing methods include noise reduction, contrast enhancement, patch generation from whole slide images (WSIs), and normalization (e.g., stain normalization in histological images) [27, 28]. However, many novel machine learning (ML)- and deep learning (DL)-based feature extractors do not require any preprocessing steps.
@@
For feature extraction, both manually designed and more recently DL-based automatic feature extractors have been utilized. Many pre-trained DL-based models, such as EfficientNet [29], and DenseNet121 [30], have been employed as feature extractors for various computer vision tasks including medical image retrieval [31, 32]. Hand-crafted or automatically extracted features of a query image are then compared to the image features in a database to find similar ones. For this comparison, various similarity measurements such as distance-based retrieval (SDR), customized query approach (CQA), or nearest-neighbor matching have been used [33]. In the context of histopathology, Yottixel [34] and FISH [35] methods select the most distinctive patches to represent WSIs through $K$-means clustering based on RGB histogram and spatial coordinate features. Yottixel employs a pre-trained DenseNet121, trained on the ImageNet dataset, to extract patch embeddings, which are then transformed into binary barcodes. FISH trains a model based on an autoencoder on the cancer genome atlas (TCGA) dataset to encode an integer index for each patch during retrieval. However, it is worth mentioning that using a feature extractor trained only on natural images based on the ImageNet dataset might be sub-optimal due to the fundamental differences between natural and medical/histopathological images.
With the advent of transformer-based models, the availability of large-scale unlabeled data and advanced vision-based or vision-language-based training methods such as contrastive language-image pre-training (CLIP) [36] or knowledge distillation with no Labels (DINO) [37], several foundation models have been developed and pre-trained on large scale data in an unsupervised or self-supervised manner. These foundation models have been used for various applications such as off-the-shelf feature extractors for computer vision tasks and include general-purpose foundation models such as OpenCLIP or medical foundation models such as UNI (for histopathological images) [23] or PanDerm (for dermatoscopic images) [38]. In this work, in addition to using well-known CNN model features for benchmarking, we employed three medical foundation models as graph node feature extractors and compared their results. The foundation models utilized include BioMedCLIP (a medical foundation model trained on diverse medical data types) and UNI and CCL (where both are medical foundation models specifically trained on histological images).
@@ 2.2 Graph Representation Learning in Histopathology
While CNNs have demonstrated impressive performance in medical image analysis, they often have trouble at capturing contextual information from neighboring areas, especially with a limited field of view. In contrast to CNNs, GNNs are better at preserving complex correlations among neighboring elements during the learning process.
In the field of computational pathology, several GNN models, such as cell graphs, tissue graphs, cell-tissue graphs, and patch graphs, have been proposed for tasks like node classification (e.g., nuclei classification) or graph classification (e.g., histologic image patch categorization) [39, 40]. However, these applications have primarily been applied in a supervised manner and often need advanced preprocessing. Using other models like HoVerNet [41] or DDU-Net [42] to generate nuclei segmentation masks as input for the graph-based model is an example of such preprocessing. As highlighted in a recent review [43], most GNN-based works in computational pathology are related to supervised classification tasks. Only a few studies have investigated histological image retrieval [24, 43, 44, 45]. However, even in these studies, the potential of attention mechanisms and adversarial training have not been explored. In our research, for the first time, we adapted and used both attention mechanism and adversarial training in variational graph autoencoders for the histological image retrieval task.
@@ 3 Methodology
The general workflow of our proposed approach for histological image retrieval is depicted in Fig. 1. In the following, we discuss the different parts of the workflow in detail.
Figure 1: The proposed general workflow of the breast histological image retrieval model involves distinct processes for training and testing. A-ARVGAE: Attention-based adversarially regularized variational graph autoencoders.
@@ 3.1 Dataset
We use two breast histopathology image datasets in this study, namely the BreakHis and the BACH dataset, which are subsequently detailed. Example images from the datasets are shown in Fig. 2.
@@
    BreakHis: This dataset contains hematoxylin and eosin (H&E)-stained breast histology microscopy images captured at various magnifications (40×40$\times$40 ×, 100×100$\times$100 ×, 200×200$\times$200 ×, 400×400$\times$400 ×) and is categorized into two primary classes, including benign and malignant classes. The dataset contains 7,909 images of which are 2,480 benign and 5,429 malignant samples with a size of $700\times 460$ pixels in Portable Network Graphics (PNG) format.
@@ 3.2 Feature Extraction
The first step in our proposed retrieval model is the feature extraction of all images in the dataset. Instead of using hand-crafted features or conventional pre-trained CNNs as feature extractors, we utilize pre-trained medical foundation models as the main feature extractors.
Unlike conventional pre-trained CNNs, the foundation models in this study are specifically trained on medical/histological images. We leverage three medical foundation models in this study, namely BioMedCLIP, UNI and CCL. BioMedCLIP is a vision-language model based on the CLIP architecture and is initially trained on the PMC15M dataset, which contains approximately 15 million image-text pairs extracted from PubMed medical articles. It utilizes a vision transformer (ViT) as its image encoder and a bidirectional encoder representation from transformers (BERT) model as its text encoder. In this study, we use only the image encoder for feature extraction. UNI is a vision model trained with a self-supervised DINO V2 [46] approach on over 100 million histological image patches sourced from 100,000 WSIs. UNI uses ViT-based models in student and teacher branches and has shown superb performance across various computer vision tasks. CCL is also a vision model, trained using a novel self-supervised, clustering-guided contrastive learning technique on over 15 million histological image patches. The CCL approach integrates two contrastive losses (a weighted InfoNCE and a group-level InfoNCE). It utilizes a clustering-guided memory bank to optimize positive and negative sample identification during contrastive learning. Further details on these models can be found in their respective studies [22, 23, 24]. We hypothesize that these models, trained on vast datasets of medical images, will yield superior features for histological image retrieval.
We also use standard baseline models as feature extractors and compare their retrieval performance with the foundation model feature extractors. For this purpose, we employ well-known pre-trained CNN models, namely VGG19 [47], MobileNetV2 [48], DenseNet121 [30], NASNetLarge [49], and EfficientNetV2M [29].